# Growth of Light-Seed Black Holes in Gas-rich Galaxies at High Redshift

### v1

@@ 3.1 No-feedback Simulation
Figure 6: Growth of MBH particle masses with feedback. Left Panel: In the left hand panel we show the growth of the active sinks as a function of time. Black and red trajectories represent BH particles that formed through supernova and direct collapse, respectively. While most of the BH particles experience negligible growth $\sim 3.5$ of all particles grow by more than 100 %. Right Panel: The Eddington ratio for all of the MBH particles that reach a final mass of greater than $10^{5}$ $\rm{M_{\odot}}$. The Eddington rates are similarly hyper-Eddington at the beginning and remain super-Eddington for up to $\sim 2\times 10^{5}$ years before galaxy disruption by feedback halts accretion.
@@
In Fig. 5, we show the final mass of the MBH particles as a function of total mass within its BHL radius ($r_{BH}$) across the entire simulation. The color bar represents the average density of the gas surrounding the MBH particles. There are two features to note; firstly we see that $1\%$ of the gas mass has been accreted onto the MBH particle on average. Secondly, the final mass is correlated with the average surrounding gas density, with the most rapidly growing MBH particles having the highest density gas surrounding them.
To attain the largest masses, the progenitor PopIII particle must remain at the center of the gas clumps they formed from. Due to the lack of feedback, there is no mechanism pushing back the gas from the PopIII particle. When the PopIII particles turn into MBH particles, the surrounding gas is therefore easily accreted - this will not always be the case in the feedback simulation (see §3.2).
There are 234 MBH particles surrounded by gas with density above $10^{-16}$ $\rm{gcm^{-3}}$ which is approaching our maximum resolution. These gas cells should convert into a PopIII particle (due to the Jeans density criteria at our resolution) but because these gas cells lie within the accretion radius of another PopIII particle, they do not convert to PopIII particles. During the stellar lifetime of the PopIII particle, these gas cells accumulate more mass and their density increases. These cells then form the basis for rapid accretion following the collapse of the PopIII particle into a MBH particle. This build-up of mass surrounding the MBH is why we see initially rapid accretion (see again the right panel of Fig 4). Once these dense gas clouds have been accreted, which occurs on a timescale of approximately $10^{5}-10^{6}$ yrs, then the growth slows dramatically with final masses of between $10^{4}$ and $10^{6}$ $\rm{M_{\odot}}$. We now examine the more realistic case of where SNe feedback can influence and in many cases suppress rapid growth.
@@ 3.2 Feedback
When SNe feedback is turned on, PopIII particles may, depending on their stellar masses, undergo SNe explosions and inject momentum (Eq. 6) to their surrounding gas. The high resolution of our simulation allows us to see the impact of individual SNe explosions. When a SNe explosion occurs, the gas surrounding the MBH particles expands spherically while transferring the momentum in radially outward directions. Recall that not all PopIII particles undergo a SNe explosion - see §2.5 for details. Out of 2181 PopIII particles, 99 experienced PISN and were deleted from the simulation.
@@
In the left panel of Fig. 6, the growth of MBH particles following their stellar phase is shown when SNe feedback is turned on. The particles that undergo a SNe explosion (and subsequently transition into a MBH particle) are shown in black. The particles which directly collapse into a MBH are shown in red. At the first glance, we can tell that just by turning SNe feedback on, the number of MBH particles that accrete over $1\ \text{M}_{\odot}$ has decreased by almost an order of magnitude compared to the no-feedback case. Only $184$ MBH particles grow by more than $1\ \text{M}_{\odot}$ compared to 1520 in the ‘no-feedback’ case. However, $16$ of these particles were still able to grow beyond $10^{5}\ \text{M}_{\odot}$. Furthermore, the feedback delays the growth as MBH particles reach $10^{5}$ $\ \text{M}_{\odot}$ in $10^{5}$ years, which is an order of magnitude longer than in the no-feedback case.
@@
Out of the 184 MBH particles that grow by more than a solar mass, 101 form through the direct collapse black hole (DCBH) channel, where there is no SNe feedback and the PopIII particle directly collapses into a MBH. As a consequence, the surrounding gas experiences no push back. These MBH particles are similar to the no-feedback simulation and their growth trajectories are shown in red in the left panel of Fig. 6. A universal trend for the growth is the appearance of flat line after a few $10^{5}$ years because the galaxy is completely disrupted by the SNe and MBH particles find efficient accretion more difficult to attain.
@@
In the right hand panel of Fig. 6, we show the Eddington ratio for each of the active sinks as a function of time. Similar to the no-feedback case, the MBH particles initially grow extremely rapidly with Eddington ratios between $10^{2}$ and $10^{4}$. We see that for MBH particles that undergo SNe, the $\rm{f_{edd}}$ begins at an earlier time compared to the ‘no-feedback’ case. Additionally, the SNe feedback decreases $\rm{f_{edd}}$ by an order of magnitude compared to the ‘no-feedback’ case. Interestingly, MBH particles that undergo SNe accrete more rapidly than those formed via the DCBH channel i.e. the rapid growth is dominated by black lines.
@@
To understand why the exploding particles grow despite SNe feedback, we show in Fig. 7 the final mass of the MBH particles as a function of the ratio between the gas momentum in the radially inward direction towards the MBH particle and momentum injected by the SNe. The gas momentum is calculated for the closest $1000$ gas cells to the progenitor PopIII particles just before the SNe explosions. The gray region is where the radially inward momentum is greater than the SNe feedback. We see that for a majority of MBH particles that grow, the radially inward momentum is larger than the SNe momentum, so feedback is inefficient in pushing back the gas. Furthermore, because the gas had expanded and re-collapsed, its angular momentum had decreased. However, this does not necessarily lead to increased accretion rates as the newly ionised gas is hotter and takes more time to collapse. Finally, in Fig. 7, we also notice a few points outside the grey area, suggesting that these MBH particles grew despite SNe successfully pushing back the surrounding gas. We find that for these particles, the shock waves produced by SNe originating from other PopIII particles had pushed gas in their direction, allowing them to quickly capture and accrete from it.
@@ 4 Discussion
The summary result from this study is that we observe multiple cases of rapid, hyper-Eddington, accretion onto light seeds buried in a dense gas-rich environment. One reason for the rapid growth we observe is that the PopIII particles do not easily decouple from the gas clumps in which they are formed in. The size of the gas clump depends on the minimum resolution of our simulations - this is a limitation of our simulations (and many others). Increasing the resolution would cause further fragmentation and formation of even smaller gas clumps. This potentially could decrease the amount of gas a MBH particle would accrete in the runaway accretion episode. A higher resolution realisation may also result in the formation of more MBH particles, dividing the available gas for accretion for an individual MBH particle.
In any case and independent of the caveats regarding resolution, the results we show here are important given the outstanding question of whether light seeds or heavy seeds are the dominant pathways to massive black holes in galaxy centres (e.g. Regan & Volonteri, 2024). It is however, not the first time the growth of light seeds has been demonstrated in numerical models (e.g. Gordon et al., 2024; Shi et al., 2023; Smith et al., 2018; Lupi et al., 2016). We now briefly review and discuss the existing literature and results and put the results from this study in context.
Recently, Gordon et al. (2024) studied the growth of light seeds in mini-halos at a higher resolution of $\Delta x\approx 10^{-4}$ pc. Using a similar sink formation and accretion prescription to us within the Enzo code, the MBH particles in their simulation similarly start accreting in high density regions, able to grow to approximately 2000-3000 $\ \text{M}_{\odot}$ within 1 Myrs. However, accretion halts because a thick disk forms around the MBH particles, which increases the gas’s angular momentum. Their MBH particles fail to grow beyond a few times $10^{3}$ $\ \text{M}_{\odot}$ for two reason: i) their higher resolution (100×100$\times$100 × our resolution) causes fragmentation to form smaller gas clouds, ii) this higher resolution also allows the formation of an accretion disk around the MBH particles which we do not see. The MBH particles are then unable to accrete gas and instead the gas accumulates on an accretion disk.
@@
The most relevant and directly comparable work to ours is that of Shi et al. (2023), who studied the growth of BH particles in large star forming molecular clouds and found hyper-Eddington growth. They found that sustained hyper-Eddington accretion occurs when the following conditions are satisfied: i) the BH particles find themselves in a dense region with gas density $\sim 100$ times that of the mean density, ii) low relative velocities between the BHs particles and the dense gas clumps (The BH particles are captured by the gas clumps). Both of these conditions are easily satisfied by the MBH particles in our simulations, which do not decouple from their formation gas clump. They also find that the surface density of the entire cloud determines the maximum mass a BH particle can grow to. In Fig. 8, we show the maximum mass of MBH particles formed as a function of surface density of the galaxy at that time. We calculate the surface density within the central $100$ pc of the galaxy. Similarly to Shi et al. (2023), we find that the final mass of the MBH particle is correlated with the surface density of the galaxy. Additionally, we observe the final mass of the MBH particles is independent of the the initial MBH particle masses (Mass immediately after the stellar lifetime ends). Another similarity arises from the dual role of feedback on accretion. While SNe feedback suppresses accretion overall, it also creates regions of colliding shock waves that prove excellent for gas capture by BH particles - again consistent with the results from Shi et al. (2023).
@@
Another study conducted by Lupi et al. (2016) found light seeds growth in the nuclear region of high-redshift massive spiral galaxies. They performed Eulerian simulations using Ramses (Teyssier, 2002) and Lagrangian simulations using Gizmo (Hopkins, 2014) at a resolution range of $0.4$ pc to $0.02$ pc (comparable if a little below ours). They show that for BH particles to grow, radiatively inefficient accretion is necessary (e.g. Madau et al., 2014). However, this is not a sufficient condition - they additionally find that the BH particle must bind to the dense gas, consistent with our results. They additionally find that the gas must be dense enough to survive the impact of feedback, again consistent with our results. As our simulations do not have radiative BH feedback, we satisfy all of the conditions they describe. Lupi et al. (2016) show extremely rapid growth of light seeds up to masses of approximately $10^{3-4}$ $\rm{M_{\odot}}~{}$depending on the different simulation setups and resolution.
@@
While the above (full hydrodynamic) studies all show rapid growth of light seeds is possible - there are clearly differences in the subgrid models employed and more importantly in the resolution in each case. Higher resolution simulations show lower overall accretion rates (e.g. Gordon et al. 2024 and Prole et al. 2022). It is therefore a caveat of our work that a higher resolution study may indeed lead to higher levels of fragmentation and less growth for our MBH particles. However, lower resolution simulations also show conflicting results and lower levels of growth. Smith et al. (2018) when analysing the Renaissance simulation suite (Xu et al., 2013, 2014; Chen et al., 2014; Ahn et al., 2015; O’Shea et al., 2015; Xu et al., 2016b, a) observe no significant growth in any of their more than 20,000 BH particles spread across a range of high redshift galaxies. The reasons for this are likely to be multiple. While their resolution is relatively low (the densities they work with are between $10^{-25}$ to $10^{-23}$ $\rm{g/cm^{3}}$, almost 7 order of magnitude lower than us), this could be expected to increase rather than suppress growth.
@@
However, even with a boost factor of $100-200$ in accretion rates, they do not get as high accretion rate as we see. Furthermore, Smith et al. (2018) observed that their BH particles did not encounter high densities regions within their cosmological halos. One of the reasons for this is due to the radiation feedback from their progenitor stars, a feature we do not model. The main difference is likely to boil down to the environments considered. The surface mass densities reached in our work are considerably higher than those seen in the Renaissance suite studied by Smith et al. (2018) and hence it may be that they either did not resolve or encounter the correct environment. Hence, while their study showed that growth is not obvious in a cosmology context, we show that given the correct conditions, rapid accretion becomes possible and BHs can grow at super (hyper)-Eddington rates. We also emphasise that our study (as many of the studies which show light seed growth) is highly idealised. A more self-consistent study is needed to allow a more robust comparison.
In future work we will explore the growth of light seed in a cosmological context including the impact of radiative feedback from PopIII particles. This additional realisation will bring our work closer to that of Renaissance but ideally at higher resolution where we can test if our results hold in more realistic settings. The additional realism of progenitor radiative feedback may contribute to decreasing the accretion rate because the gas clump may be disrupted by this feedback even before the SNe feedback sets in. Similarly, we also neglect feedback from BH accretion which will be included in future work.

### v2

@@ 3.1 No-feedback Simulation
Figure 9: Growth of MBH particle masses with feedback. Left Panel: In the left hand panel we show the growth of the active sinks as a function of time. Blue and red trajectories represent BH particles that formed through supernova and direct collapse, respectively. While most of the BH particles experience negligible growth $\sim 3.5$ of all particles grow by more than 100 %. Right Panel: The Eddington ratio for all of the MBH particles that reach a final mass of greater than $10^{5}$ $\rm{M_{\odot}}$. The Eddington rates are similarly hyper-Eddington at the beginning and remain super-Eddington for up to $\sim 2\times 10^{5}$ yr before galaxy disruption by feedback halts accretion.
@@
In Fig. 7, we show the final mass of the MBH particles as a function of total mass within its BHL radius ($r_{BH}$) throughout the simulation. The colourbar represents the average density of the gas surrounding the MBH particles. There are two features to note; first, we see that $1\%$ of the gas mass has been accreted onto the MBH particle on average. Secondly, the final mass is correlated with the average surrounding gas density, with the most rapidly growing MBH particles having the highest density gas surrounding them.
To attain the largest masses, the progenitor PopIII particle must remain at the centre of the gas clumps they formed from. Due to the lack of feedback, there is no mechanism that pushes the gas back from the PopIII particle. When PopIII particles turn into MBH particles, the surrounding gas is therefore easily accreted; this will not always be the case in the feedback simulation (see §3.2). The presence of radiative feedback from PopIII particles would also push back the gas, which added with supernova feedback may decrease the number of growing BHs (see §5).
There are 234 MBH particles surrounded by gas with density greater than $10^{-16}$ $\rm{g/cm^{3}}$, which is approaching our maximum resolution. These gas cells should convert into a PopIII particle (due to the Jeans density criteria at our resolution), but because these gas cells lie within the accretion radius of another PopIII particle, they do not convert to PopIII particles. During the stellar lifetime of the PopIII particle, these gas cells accumulate more mass and their density increases. These cells then form the basis for rapid accretion following the collapse of the PopIII particle into a MBH particle. This build-up of mass surrounding the MBH is why we see initially rapid accretion (see again the right panel of Fig 6). Once these dense gas clouds have been accreted, which occurs on a timescale of approximately $10^{5}-10^{6}$ yr, then the growth slows dramatically with final masses between $10^{4}$ and $10^{6}$ $\rm{M_{\odot}}$. We now examine the more realistic case where SNe feedback can influence and in many cases suppress rapid growth.
@@ 3.2 Feedback
When the SNe feedback is turned on, PopIII particles may, depending on their stellar masses, undergo SNe explosions and inject momentum (Eq. 6) into their surrounding gas. To see the impact of supernova feedback, we show face-on gas density projection plots in Fig. 8. We show three instances of the simulation at $t=382,577,773$ kyr for the entire galaxy (top panel) and zoom in on the densest clump in the IC (bottom panel). The high resolution of our simulation allows us to see the impact of individual SNe explosions. When a SNe explosion occurs, the gas surrounding the MBH particles expands spherically while transferring the momentum in radially outward directions. This can be seen in the top-right panel at $t=773$ kyr where a large supernova occurs at the top side of the galaxy. Due to multiple supernovae occurring in a short time period, the gas is injected with a lot of energy, which completely disrupts the galaxy. This can be seen in the bottom panels, where the dense clumps and filaments are completely disrupted by the supernovae. Recall that not all PopIII particles undergo a SNe explosion - see §2.5 for details. Out of 2181 PopIII particles, 99 experienced PISN and were deleted from the simulation.
@@
In the left panel of Fig. 9, the growth of MBH particles following their stellar phase is shown when the SNe feedback is turned on. The particles that undergo a SNe explosion (and subsequently transition into a MBH particle) are shown in blue. The particles that directly collapse into a MBH are shown in red. At first glance, we can tell that just by turning on SNe feedback, the number of MBH particles that accrete over $1\ \text{M}_{\odot}$ has decreased by almost an order of magnitude compared to the no-feedback case. Only $184$ MBH particles grow by more than $1\ \text{M}_{\odot}$ compared to 1520 in the no-feedback case. However, $16$ of these particles were still able to grow beyond $10^{5}\ \text{M}_{\odot}$. Furthermore, feedback delays growth as MBH particles reach $10^{5}$ $\ \text{M}_{\odot}$ in $10^{5}$ yr, which is an order of magnitude longer than in the no-feedback case.
@@
Out of the 184 MBH particles that grow by more than a solar mass, 101 form through the direct collapse black hole (DCBH) channel, where there is no SNe feedback and the PopIII particle directly collapses into a MBH. As a consequence, the surrounding gas does not experience any push back. These MBH particles are similar to the no-feedback simulation, and their growth trajectories are shown in red in the left panel of Fig. 9. A universal trend for the growth is the appearance of flat line after a few $10^{5}$ yr because the galaxy is completely disrupted by the SNe and MBH particles find efficient accretion more difficult to attain.
@@
In the right panel of Fig. 9, we show the Eddington ratio for each of the active sinks as a function of time. Similarly to the no-feedback case, MBH particles initially grow extremely rapidly with Eddington ratios between $10^{2}$ and $10^{4}$. We see that for MBH particles that undergo SNe, the $\rm{f_{edd}}$ begins at an earlier time compared to the no-feedback case. Furthermore, SNe feedback decreases $\rm{f_{edd}}$ by an order of magnitude compared to the no-feedback case. Interestingly, MBH particles that undergo SNe accrete more rapidly than those formed via the DCBH channel, i.e. the rapid growth is dominated by blue lines.
@@
To understand why the exploding particles grow despite SNe feedback, we show in Fig. 10 the final mass of the MBH particles as a function of the ratio between the gas momentum in the radially inward direction towards the MBH particle and the momentum injected by the SNe. The gas momentum is calculated for the closest $1000$ gas cells to the progenitor PopIII particles just before the SNe explosions. The grey region is where the radially inward momentum is greater than the SNe feedback. We see that for a majority of MBH particles that grow, the radially inward momentum is larger than the SNe momentum, so feedback is inefficient in pushing back the gas. Furthermore, because the gas had expanded and re collapsed, its angular momentum had decreased. However, this does not necessarily lead to increased accretion rates as the newly ionised gas is hotter and takes more time to collapse. Finally, in Fig. 10, we also notice a few points outside the grey area, suggesting that these MBH particles grew despite SNe successfully pushing back the surrounding gas. We find that for these particles, the shock waves produced by SNe originating from other PopIII particles had pushed gas in their direction, allowing them to quickly capture and accrete from it.
@@ 4 Discussion
The summary result from this study is that we observe multiple cases of rapid, hyper-Eddington, accretion onto light seeds buried in a dense gas-rich environment. One reason for the rapid growth we observe is that the PopIII particles do not easily decouple from the gas clumps in which they are formed. The size of the gas clump depends on the minimum resolution of our simulations - this is a limitation of our simulations (and many others). Increasing the resolution would cause further fragmentation and formation of even smaller gas clumps. This potentially could decrease the amount of gas a MBH particle would accrete in the runaway accretion episode. A higher resolution realisation may also result in the formation of more MBH particles, dividing the gas available for accretion by an individual MBH particle.
In any case and independent of the caveats regarding resolution, the results we show here are important given the outstanding question of whether light seeds or heavy seeds are the dominant pathways to massive black holes in galaxy centres (e.g. Regan & Volonteri, 2024). It is however, not the first time that the growth of light seeds has been demonstrated in numerical models (e.g. Gordon et al., 2024; Shi et al., 2023; Smith et al., 2018; Lupi et al., 2016). We now briefly review and discuss the existing literature and results and put the results from this study in context.
Recently, Gordon et al. (2024) studied the growth of light seeds in mini-halos at a higher resolution of $\Delta x\approx 10^{-4}$ pc. Using a similar sink formation and accretion prescription to us within the Enzo code, the MBH particles in their simulation similarly start accreting in high density regions, able to grow to approximately 2000-3000 $\ \text{M}_{\odot}$ within 1 Myr. However, accretion halts because a thick disk forms around the MBH particles, which increases the gas’s angular momentum. Their MBH particles fail to grow beyond a few times $10^{3}$ $\ \text{M}_{\odot}$ for two reasons: i) their higher resolution (100×100$\times$100 × our resolution) causes fragmentation to form smaller gas clouds, and ii) this higher resolution also allows the formation of an accretion disk around the MBH particles which we do not see. The MBH particles are then unable to accrete gas and instead the gas accumulates on an accretion disk.
@@
The most relevant and directly comparable work to ours is that of Shi et al. (2023), who studied the growth of BH particles in large star forming molecular clouds and found hyper-Eddington growth. They found that sustained hyper-Eddington accretion occurs when the following conditions are satisfied: i) the BH particles find themselves in a dense region with gas density $\sim 100$ times that of the mean density, ii) low relative velocities between the BHs particles and the dense gas clumps (The BH particles are captured by the gas clumps). Both of these conditions are easily satisfied by the MBH particles in our simulations, which do not decouple from their formation gas clump. They also find that the surface density of the entire cloud determines the maximum mass a BH particle can grow to. In Fig. 11, we show the maximum mass of MBH particles formed as a function of surface density of the galaxy at that time. We calculate the surface density within the central $100$ pc of the galaxy. Similarly to Shi et al. (2023), we find that the final mass of the MBH particle is correlated with the surface density of the galaxy. Additionally, we observe that the final mass of the MBH particles is independent of the initial MBH particle masses (mass immediately after the stellar lifetime ends). Another similarity arises from the dual role of feedback on accretion. While SNe feedback suppresses accretion overall, it also creates regions of colliding shock waves that prove excellent for gas capture by BH particles - again consistent with the results from Shi et al. (2023).
@@
Another study conducted by Lupi et al. (2016) found light seeds growth in the nuclear region of high-redshift massive spiral galaxies. They performed Eulerain simulations using Ramses (Teyssier, 2002) and Lagrangian simulations using Gizmo (Hopkins, 2014) at a resolution range of $0.4$ pc to $0.02$ pc (comparable if a little below ours). They show that for BH particles to grow, radiatively inefficient accretion is necessary (e.g. Madau et al., 2014). However, this is not a sufficient condition - they additionally find that the BH particle must bind to the dense gas, consistent with our results. They additionally found that the gas must be dense enough to survive the impact of feedback, again consistent with our results. As our simulations do not have radiative BH feedback, we satisfy all of the conditions described. Lupi et al. (2016) show extremely rapid growth of light seeds up to masses of approximately $10^{3-4}$ $\rm{M_{\odot}}~{}$depending on the different simulation setups and resolution.
@@
While the above (full hydrodynamic) studies all show rapid growth of light seeds is possible - there are clearly differences in the subgrid models employed and more importantly in the resolution in each case. Higher resolution simulations show lower overall accretion rates (e.g. Gordon et al. 2024 and Prole et al. 2022). It is therefore a caveat of our work that a higher resolution study may indeed lead to higher levels of fragmentation and less growth for our MBH particles. However, lower resolution simulations also show conflicting results and lower levels of growth. Smith et al. (2018) when analysing the Renaissance simulation suite (Xu et al., 2013, 2014; Chen et al., 2014; Ahn et al., 2015; O’Shea et al., 2015; Xu et al., 2016b, a) observes no significant growth in any of their more than 20,000 BH particles spread across a range of high redshift galaxies. The reasons for this are likely to be multiple. While their resolution is relatively low (the densities they work with are between $10^{-25}$ to $10^{-23}$ $\rm{g/cm^{3}}$, almost 7 orders of magnitude lower than ours), this could be expected to increase rather than suppress growth.
@@
However, even with a boost factor of $100-200$ in accretion rates, they do not get as high accretion rate as we see. Furthermore, Smith et al. (2018) observed that their BH particles did not encounter high densities regions within their cosmological halos. One of the reasons for this is due to the radiation feedback from their progenitor stars, a feature we do not model. The main difference is likely to boil down to the environments considered. The surface mass densities reached in our work are considerably higher than those seen in the Renaissance suite studied by Smith et al. (2018), and hence it may be that they either did not resolve or encounter the correct environment. Hence, while their study showed that growth is not obvious in a cosmology context, we show that given the correct conditions, rapid accretion becomes possible and BHs can grow at super (hyper)-Eddington rates. We also emphasise that our study (as many of the studies which show light seed growth) is highly idealised. A more self-consistent study is needed to allow a more robust comparison.
5 Caveats and Future Work
Our study, given its idealised setup, has a number of caveats, which we now discuss. The first caveat of our study is the short timescale on which we run the simulations ($\sim 10$ Myr). As a consequence, we shorten our stellar lifetimes by a factor of 100 so that the remnant BH particles form early. This leads to the majority of PopIII particles having lifetimes shorter than $10^{5}$ yr. This in turn means that PopIII stars have very little time to decouple from their formation gas clumps. Is it possible that longer (more realistic) lifetimes may result in the PopIII stars decoupling from their formation cloud and subsequently the remnant BH being born into lower density environments. This would lead to less BHs reaching $10^{5}\ \text{M}_{\odot}$. Additionally, in the feedback simulations, because of shorter stellar lifetimes, we have multiple supernovas going off within a short time, leading to complete disruption of the galaxy. This, of course, has the opposite effect and potentially suppresses growth.
Another caveat of our work is the absence of radiative feedback from PopIII particles. Including radiative feedback from stellar populations would bring our work closer to that of Renaissance (e.g. Smith et al., 2018). We plan to do this in a forthcoming study. The additional realism of progenitor radiative feedback may contribute to decreasing the accretion rate because the gas clumps may be disrupted by this feedback even before the SNe feedback sets in. However, this is not a certainty. High resolution simulations of the effects of radiative feedback from PopIII stars have shown formation of small H II regions around PopIII stars, with the radiation remaining trapped inside dense gas clumps (Jaura et al., 2022).
In addition to neglecting feedback from the stellar populations, we neglect feedback from the accreting black holes. This again would work to suppress accretion and is a limitation of our work here (and may set an upper limit to our results).
In future work, we will explore the growth of light seed in a cosmological context, including the impact of radiative feedback from PopIII particles and from the accreting BHs. By expanding our dataset to a cosmological setting, we plan to explore the conditions required for light seed growth in a more self-consistent and less idealised fashion.